# Sparsity May Be All You Need: Sparse Random Parameter Adaptation

### v2

@@ Abstract
Full fine-tuning of large language models for alignment and task adaptation has become prohibitively expensive as models have grown in size. Parameter-Efficient Fine-Tuning (PEFT) methods aim at significantly reducing the computational and memory resources needed for fine-tuning these models by only training on a small number of parameters instead of all model parameters. Currently, the most popular PEFT method is the Low-Rank Adaptation (LoRA), which freezes the parameters of the model to be fine-tuned and introduces a small set of trainable parameters in the form of low-rank matrices. We propose simply reducing the number of trainable parameters by randomly selecting a small proportion of the model parameters to train on. In this paper, we compare the efficiency and performance of our proposed approach with PEFT methods, including LoRA, as well as full parameter fine-tuning.
@@ 1 Introduction
Figure 1: The proposed SpaRTA method which randomly chooses a subset of parameter indices vector $\phi$ from pre-trained model parameters $\theta_{\text{PT}}$ and updates them via adapter $\Delta_{\phi}$.
Low Rank Adaptation (LoRA), first introduced by Hu et al. (2022), currently remains the most popular PEFT technique. LoRA freezes all the pre-trained model parameters $\theta_{\text{PT}}$ and introduces trainable low-rank matrices (e.g., $B$, $A$) for a pre-selected subset of matrix tensors to represent the changes ($\Delta=BA$) needed for adapting the model to a new task. The adapted model parameters are given by $\theta_{\text{PT}}+\Delta$. Memory and computational efficiency are achieved by optimizing only over the parameters of these newly added, but significantly smaller, matrices.
@@
In this work, we propose a different approach where $\Delta$ is not factorized into low rank matrices but rather chosen to be a random subset of the model parameters. This Sparse Random parameTer Adaptation (SpaRTA) method imposes a sparsity constraint on the adapter that can be easily controlled. By changing the desired sparsity, one can change the number of adaptation parameters. Regardless of how the selected parameters are sampled from the model parameters, subsequent updates only affect these parameters. This sparsity constraint and randomness of selected parameters is in contrast to techniques such as LoRA that effectively affect all parameters in $\theta_{\text{PT}}$. See Figure 1 for an illustration of the method. Generally, one samples $m$ parameters from the pre-trained model $\theta_{\text{PT}}$, stores the indices of these parameters in $\phi$, and uses adapter $\Delta_{\phi}$ to fine-tune $\theta_{\text{PT}}$.
@@ 2 Motivation
This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters: SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero). This $\Delta$-sparsity also helps in reducing overfitting, as pre-trained models typically have more than enough capacity to learn the often limited amount of (labeled) data used for fine-tuning. There is no guarantee of $\Delta$-sparsity in LoRA, but this is a desired property since it reduces parameter interference (Yadav et al., 2023) when merging fine-tuned models.
SpaRTA, like LoRA, reduces the number of gradient computations during training (when compared with full Fine-Tuning), and ultimately has the same inference cost as the original model (after merging back the sparse $\Delta$ into the model). The properties of low memory, low computation needs in training, identical inference time as the original model are all quite desirable for an adaptation technique. SpaRTA has all these properties plus the unique added benefit of producing only sparse changes in a small number of the parameters of the original model that can be beneficial for merging multiple SpaRTA adapters.
@@ 2.1 Is low-rank adaption necessary?
In Appendix A, we show that the changes (i.e., $\Delta$) in weight matrix during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b and mistral-7b. This indicates that LoRA works, not particularly because of its low-rank constraint, but rather due to the capacity reduction achieved since LoRA fine-tuning with is typically done on limited training data. Such insight also hints that any constraint which reduces the capacity of the original model could perform competitively, motivating our approach that selects a small number of parameters of the original model to be updated during training.
@@ 3 Related work
The last few years has seen many advances in fine-tuning Large Language Models (LLMs). Perhaps the most well-known and used method in practice is LoRA Hu et al. (2022), which has spurred many variants including: DoRA, which adapts only the directions of pre-trained weights Liu et al. (2024); VeRA, which shares low-rank matrices across layers Kopiczko et al. (2024); AdaLoRA, which adaptively allocates parameter budgets among weight matrices according to their importance Zhang et al. (2023); and SoRA, which dynamically adjusts the intrinsic rank during adaptation Ding et al. (2023). Each of these methods has a different structure to the $\Delta$ being optimized, with the commonality being the training of some function of low-rank matrices. AdaLoRa and SoRA are also dynamic methods, which adjust the function during adaptation.
@@ 4 SpaRTA: Adapting a Random Subset of Model Parameters
Suppose the vectorized parameters of the pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) of the model is performed with a labeled dataset characterizing a task. FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm, e.g., Adam from Kingma and Ba (2015), to maximize the conditional probability of the labels in the training dataset under the model. The FT model is then given by $\theta_{\text{FT}}=\theta_{\text{PT}}+\Delta_{\text{FT}}$ where $\Delta_{\text{FT}}\in\mathbb{R}^{n}$.
@@
SpaRTA proposes to randomly select a small subset of the model parameters to optimize while freezing the rest. The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of parameters selected. This allows our approach to have a drastically lower memory footprint than FT by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute.
In our SpaRTA’s implementation [^1] https://github.com/IBM/SpaRTA, we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$. The pseudocode for SpaRTA is given in Algorithm 1.
@@ 5. Update $\Delta_{\phi}$ with the Adam Optimizer
Generating the index set $\phi$ in SpaRTA is done by sampling from a Bernoulli independently and including each scalar parameter in $\phi$ with probability $m/n$. Hence, SpaRTA uses $m$ trainable parameters out of $n$ in expectation. Finally, inference is performed, similarly to LoRA, by merging the fine-tuned $\Delta$ into $\theta_{\text{PT}}$ and then making the necessary forward passes. Thus, SpaRTA does not introduce any additional inference latency when deploying the resulting fine-tuned model.
@@ 5 Memory Usage
Recall that SpaRTA freezes $n-m$ ($m\ll n$) of the model parameters. We define sparsity as $s=1-m/n\in(0,1)$, the percentage of frozen model parameters (e.g., if $1\%$ of parameters are trainable, then the sparsity is $99\%$). Subsequently, density is defined as $k=m/n=1-s$, the percentage of trainable model parameters. In practice, for a chosen sparsity $s$, one can freeze a model parameter with probability $s$, expecting a total sparsity percentage of $s$ over all model parameters. Thus, in expectation, $k=m/n$ percent of the model parameters are chosen as trainable, for a total of $n\,k=m$ trainable parameters.
For SpaRTA, only $\Delta_{\phi}$ (of size $m$) is trainable, which is significantly smaller than the total number of trainable parameters for FT since $m\ll n$. However, the indices of these randomly chosen model parameters must be recorded into $\phi$, adding to the memory requirements. Indices can be stored in $16$-bit integers for all the Transformer models (Vaswani et al., 2017) considered in this paper.
@@
The values in $\Delta_{\phi}$ are of the same type as the model parameters, e.g., using $16$-bit brain floating-point (bfloat16) tensors. Thus, SpaRTA requires up to $m\,(2\times 16+16)=3m\times 16$ bits of extra memory to specify the index $\phi$ and delta $\Delta_{\phi}$ tensors. That is $3\,k$ times more memory than the original model, which requires just $n\times 16$ bits for storing its parameters. For instance, using SpaRTA on a model with sparsity 80%, 90%, 95% and 99% would require up to 60%, 30%, 15% and 3% more memory, respectively.
We next analyze [^2] This analysis does not include memory requirements associated with model buffers (e.g., those used to track running statistics for layer normalization) because they are relatively small and the same for both the Full FT and SpaRTA. SpaRTA’s memory savings during training. Optimizing the full set of model parameters (FT) using Adam (Kingma and Ba, 2015) requires memory for the parameters, their gradients with their (adaptive) first and second moments estimates. This requires $4\,n\times 16$ bits of memory when using a bfloat16 representation.
In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits for the statistics needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits). Memory savings appear then if and only if
$m\,(4\!\times\!16+2\!\times\!16)+n\times 16$ $<4n\!\times\!16,$ (1)
$kn\,(6\!\times\!16)$ $<3n\!\times\!16,$
$k$ $<\frac{1}{2},$
that is, SpaRTA is a real PEFT method iff $k<0.5$, i.e., a sparsity higher than 50% is required ($s\!>\!0.5$), making less than 50% of the model parameters trainable. For instance, using SpaRTA on a model with sparsity $s=$ 80%, 90%, 95%, and 99% requires 45%, 60%, 67.5% and 73.5% less memory than full parameter FT, respectively, see Table 1. Additional savings are possible regarding storing indices given a fixed random number generator; in this case, the random path of selected parameters to train can be derived given a fixed seed.
@@
[…]
@@ 6 Experimental Setup
We next detail our experimental framework. Motivation is first given for the tasks followed by a description of models to be used and how these models are adapted (with manipulations) to the tasks.
@@ 6.2 Language Models
Starting with available open-weight language models, the goal is to adapt them to perform a selected NLU task. Two types of trained models are used: base and instruction-tuned models, where the latter have additionally been trained to follow users’ intent when prompted, such as answering questions or responding to instructions. Specifically, we consider the following language models: gemma-2b and gemma-2b-it from the Gemma family (Team et al., 2024), and mistral-7b and mistral-7b-it from the Mistral [^3] We use v0.3 for both models. family (Jiang et al., 2023). The "it" suffix refers to instruction-tuned models. They are of particular interest for our experiments as they will show results for models with two different numbers of parameters. All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face [^4] https://huggingface.co/. Note that when using an instruction-following (it) model, inputs are formatted according to the conventions established when training the model on instructions.
@@ 7.1 IMDB
Table 3: Full table of loss and accuracy on the SST-2 task after applying our SpaRTA approach to different types of parameters in the Gemma 2B model, given that the same number of trainable parameters are selected. Results are averaged across 10 random seeds
gemma-2b gemma-2b-it mistral-7b mistral-7b-it method density (%) loss acc. loss acc. loss acc. loss acc. PT (zero-shot) 0.425 92.0% 0.435 86.1% Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.050.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1%
Table 4: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods. We also report zero-shot performance of instruction following PT models. For training details see Table 16.
Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. avg. PT (zero-shot) 1.34 40.5 1.77 43.7 0.75 60.7 1.41 48.3 2.91 30.9 0.91 64.6 0.85 0.7 0.88 68.5 44.7 SpaRTA (5%) ALL 0.17 93.7 0.54 81.0 0.14 95.0 0.26 89.0 0.33 87.2 0.35 85.3 0.39 56.2 0.36 84.4 84.0 SpaRTA (0.5%) 0.16 94.0 0.43 80.7 0.15 94.7 0.24 89.7 0.32 87.8 0.34 86.5 0.45 54.9 0.37 84.7 84.1 SpaRTA (0.037%) 0.18 93.1 0.48 76.9 0.16 94.5 0.26 88.8 0.35 86.4 0.36 84.5 0.45 54.8 0.40 83.6 82.8 LoRA (0.037%) Q,V 0.18 93.4 0.44 78.7 0.14 95.1 0.26 89.1 0.33 87.8 0.33 85.4 0.41 55.0 0.36 84.0 83.6 DoRA (0.037%) 0.18 93.4 0.45 78.0 0.15 95.6 0.26 89.0 0.33 87.8 0.33 85.5 0.41 57.1 0.36 84.4 83.8 SpaRTA (0.037%) 0.19 92.8 0.46 77.6 0.16 94.8 0.26 88.6 0.36 86.3 0.36 84.2 0.45 54.5 0.39 83.4 82.8 SpaRTA (0.037%) O,V 0.17 93.5 0.44 80.7 0.15 94.9 0.26 88.9 0.34 87.2 0.35 85.7 0.42 55.9 0.37 85.0 84.0
Table 5: Test loss and accuracy of model gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods. For training details see Appendix E.2.
Table 4 shows the results for IMDB, where each model is asked to classify a review as positive or negative. Each model is fine-tuned using the following adaptation methods: (i) Full parameter fine-tuning (Full FT) where all model parameters are optimized; (ii) SpaRTA for different density levels 5%, 0.5%, 0.05%, with the last allowing SpaRTA to have approximately the same number of trainable parameters as LoRA; (iii) LoRA with rank $r=8$ equivalent to about $0.05\%$ of trainable parameters compared to the model full parameter size.

### v3

@@ Abstract
Full fine-tuning of large language models for alignment and task adaptation has become prohibitively expensive as models have grown in size. Parameter-Efficient Fine-Tuning (PEFT) methods aim at significantly reducing the computational and memory resources needed for fine-tuning these models by only training on a small number of parameters instead of all model parameters. Currently, the most popular PEFT method is the Low-Rank Adaptation (LoRA), which freezes the parameters of the model and introduces a small set of trainable parameters in the form of low-rank matrices. We propose simply reducing the number of trainable parameters by randomly selecting a small proportion of the model parameters to train on, while fixing all other parameters, without any additional prior assumptions such as low-rank structures. In this paper, we compare the efficiency and performance of our proposed approach to other PEFT methods as well as full parameter fine-tuning. We find our method to be competitive with LoRA when using a similar number of trainable parameters. Our findings suggest that what truly matters for a PEFT technique to perform well is not necessarily the specific adapter structure, but rather the number of trainable parameters being used.
@@ 1 Introduction
Figure 1: The proposed SpaRTA method which randomly chooses a small subset of parameters from $\theta_{\text{PT}}$, stores the indices of the selected parameters in $\phi$ and updates the model via adapter $\Delta_{\phi}$.
Low Rank Adaptation (LoRA), first introduced by Hu et al. (2022), currently remains the most popular PEFT technique. LoRA freezes all the pre-trained model parameters $\theta_{\text{PT}}$ and introduces trainable low-rank matrices (e.g., $B$, AA) to represent the changes ($\Delta=BA$) needed for adapting the model to a new task. The adapted model parameters are given by $\theta_{\text{PT}}+\Delta$. Memory and computational efficiency are achieved by optimizing only over the parameters of these newly added, but significantly smaller, matrices.
@@
In this work, we propose a different approach where $\Delta$ is not factorized into low rank matrices but rather chosen to be a random subset of the model parameters. This Sparse Random parameTer Adaptation (SpaRTA) method imposes a sparsity constraint on the adapter that can be easily controlled. By changing the desired sparsity, one can change the number of adaptation parameters. Regardless of how the selected parameters are sampled from the model parameters, subsequent updates only affect these parameters. This sparsity constraint and randomness of selected parameters is in contrast to techniques such as LoRA that effectively affect all parameters in $\theta_{\text{PT}}$. See Figure 1 for an illustration of the method. Generally, one samples $m$ parameters from the pre-trained model $\theta_{\text{PT}}$, stores their indices in $\phi$, and uses adapter $\Delta_{\phi}$ to fine-tune $\theta_{\text{PT}}$.
@@ 2 Motivation
This motivates our approach, SpaRTA, which produces a performant fine-tuned model by directly adapting only a small percentage of the pre-trained model parameters. SpaRTA randomly selects the (scalar) parameters to train and freezes the remaining parameters (i.e., setting the corresponding $\Delta$ values to zero). This $\Delta$-sparsity also helps in reducing overfitting, as pre-trained models typically have more than enough capacity to learn the often limited amount of (labeled) data used for fine-tuning. There is no guarantee of $\Delta$-sparsity in LoRA, but this is a desired property since it reduces parameter interference (Yadav et al., 2023) when merging fine-tuned models.
SpaRTA, like LoRA, reduces the number of gradient computations during training (when compared with full fine-tuning), and ultimately has the same inference cost as the original model (after merging back the sparse $\Delta$ into the model). For an adaptation technique, having lower memory and computation needs during training as well as identical inference time to the original model are all quite desirable properties. SpaRTA has all these properties plus the unique added benefit of producing only sparse changes in a small number of the parameters of the original model that can be beneficial for merging multiple SpaRTA adapters.
@@ 2.1 Is Low-Rank Adaption Necessary?
In Appendix A, we show that the changes (i.e., $\Delta$) in weight matrix during full parameter fine-tuning are, in fact, not generally low-rank for capable models such as gemma-2b-it and mistral-7b-it. This indicates that LoRA works, not particularly because of its low-rank constraint, but rather due to the reduction in model capacity achieved by LoRA when fine-tuning on limited training data, as it is typically done in task adaption. Such insight also hints that any constraint reducing the capacity of the original model could perform competitively, motivating our approach which selects a small number of parameters from the original model to be updated during training.
@@ 3 Related Work
The last few years has seen many advances in PEFT methods. Perhaps the most well-known and used method in practice is LoRA Hu et al. (2022), which has spurred many variants including: DoRA, which adapts only the directions of pre-trained weights Liu et al. (2024); VeRA, which shares low-rank matrices across layers Kopiczko et al. (2024); AdaLoRA, which adaptively allocates parameter budgets among weight matrices according to their importance Zhang et al. (2023); and SoRA, which dynamically adjusts the intrinsic rank during adaptation Ding et al. (2023). Each of these methods has a different structure to the $\Delta$ being optimized, with the commonality being the training of some function of low-rank matrices. AdaLoRA and SoRA are also dynamic methods, which adjust the function during adaptation.
@@ 4 SpaRTA: Adapting a Random Subset of Model Parameters
Suppose the parameters of a pre-trained language model are $\theta_{\text{PT}}\in\mathbb{R}^{n}$ where $n$ is the number of parameters, and full parameter fine-tuning (FT) is performed with a labeled dataset characterizing a task. FT typically updates $\theta_{\text{PT}}$ using a stochastic first-order gradient-based optimization algorithm, e.g., Adam from Kingma and Ba (2015), to maximize the conditional probability of the labels in the training dataset under the model. The FT model is then given by $\theta_{\text{FT}}=\theta_{\text{PT}}+\Delta_{\text{FT}}$ where $\Delta_{\text{FT}}\in\mathbb{R}^{n}$.
@@
SpaRTA proposes to randomly select a small subset of the model parameters to optimize while freezing the rest. The model parameters $\theta\in\mathbb{R}^{n}$ are partitioned into trainable $\theta_{\text{T}}\in\mathbb{R}^{m}$ and frozen $\theta_{\text{F}}\in\mathbb{R}^{n-m}$ ones, with $m\ll n$ being the number of selected parameters. This allows our approach to have a drastically lower memory footprint than FT by reducing the size of the optimizer state as well as faster training by reducing the number of gradients to compute. This is similar with respect to LoRA in that both reduce the memory footprint by optimizing a small set of parameters (the adapter) while freezing most (SpaRTA) or all (LoRA) of the model parameters.
In our SpaRTA implementation [^1] Code is available at https://github.com/IBM/SpaRTA., we introduce: (i) non-trainable indices $\phi\in\mathbb{R}^{m}$ containing the indices of the randomly selected elements in $\theta_{\text{PT}}$, and (ii) trainable parameters $\Delta_{\phi}\in\mathbb{R}^{m}$ representing the subset of $\Delta$ that SpaRTA learns at indices $\phi$. The pseudocode for SpaRTA is given in Algorithm 1.
@@ 5. Update $\Delta_{\phi}$ with the Adam Optimizer
Generating the index set $\phi$ in SpaRTA is done by sampling from a Bernoulli independently and including each scalar parameter in $\phi$ with probability $m/n$. Hence, SpaRTA uses $m$ trainable parameters out of $n$ in expectation. Finally, inference is performed, similarly to LoRA, by merging the fine-tuned $\Delta$ into $\theta_{\text{PT}}$ and then making the necessary forward passes. Thus, SpaRTA does not introduce any additional inference latency.
@@ 5 Memory Usage
Recall that SpaRTA freezes nmn-m ($m\ll n$) of the model parameters. We define sparsity as $s=1-m/n\in(0,1)$, the percentage of frozen model parameters (e.g., if $1\%$ of parameters are trainable, then the sparsity is $99\%$). Subsequently, density is defined as k=m/n=1sk=m/n=1-s, the percentage of trainable model parameters. In practice, for a chosen sparsity $s$, one can freeze a model parameter with probability $s$, expecting a total sparsity percentage of $s$ over all model parameters. Thus, in expectation, $k=m/n$ percent of the model parameters are chosen as trainable, for a total of nk=mn\,k=m trainable parameters.
For SpaRTA, only $\Delta_{\phi}$ (of size $m$) is trainable, which is significantly smaller than the total number of model parameters since $m\ll n$. However, the indices of these randomly chosen parameters must be recorded into $\phi$, adding to the memory requirements. Indices can be stored in $16$-bit integers for all the Transformer models (Vaswani et al., 2017) considered in this paper.
@@
The values in $\Delta_{\phi}$ are of the same type as the model parameters, e.g., using $16$-bit brain floating-point (bfloat16) tensors. Thus, SpaRTA requires up to $m\,(2\times 16+16)=3m\times 16$ bits of extra memory to specify the index $\phi$ and delta $\Delta_{\phi}$ tensors. That is 3k3\,k times more memory than the original model, which requires just $n\times 16$ bits for storing its parameters. For instance, using SpaRTA on a model with sparsity 80%, 90%, 95% and 99% would require up to 60%, 30%, 15% and 3% more memory, respectively.
We next analyze [^2] This analysis does not include memory requirements associated with model buffers (e.g., those used to track running statistics for layer normalization) because they are relatively small and the same for both the Full FT and SpaRTA. SpaRTA’s memory savings during training. Optimizing the full set of model parameters (FT) using Adam (Kingma and Ba, 2015) requires memory for the parameters, their gradients, and their (adaptive) first and second moments estimates. This requires $4\,n\times 16$ bits of memory when using a bfloat16 representation.
In contrast, SpaRTA only optimizes $\Delta_{\phi}$, requiring a total of $m\,(4\times 16+2\times 16)+n\times 16$ bits of memory: (i) $m\,(4\times 16)$ bits needed by Adam to optimize $\Delta_{\phi}$ of size $m$, (ii) $m\,(2\times 16)$ bits for $\phi$ with the indices identifying the model parameters associated with $\Delta_{\phi}$, and (iii) the PT model parameters ($n\times 16$ bits). Memory savings appear then if and only if
$m\,(4\!\times\!16+2\!\times\!16)+n\times 16$ $<4n\!\times\!16$ (1)
$kn\,(6\!\times\!16)$ $<3n\!\times\!16$
$k$ $<0.5,$
that is, SpaRTA is a real PEFT method iff $k<0.5$, i.e., a sparsity higher than 50% is required (s>0.5s\!>\!0.5), making less than 50% of the model parameters trainable. For instance, using SpaRTA on a model with sparsity $s=$ 80%, 90%, 95%, and 99% requires 45%, 60%, 67.5% and 73.5% less memory than full parameter FT, respectively, see Table 1. Additional savings are possible regarding storing indices given a fixed random number generator; in this case, the random path of selected parameters to train can be derived given a fixed seed.
@@
Table 1: SpaRTA memory usage efficiency during training for $n=\text{2B}$ and 7B parameter models. Memory in Gigabytes (GB) for storing the parameters on disk (storage) as well as for FT with Adam on (i) the full set of parameters (Full FT) equivalent to 0% sparsity (ii) a sparse random subset (SpaRTA) for different sparsity percentages from 50% to 99%.
@@ 6 Experimental Setup
We next detail our experimental framework. Motivation is first given for the tasks followed by a description of models to be used and how they are adapted (with manipulations) to the tasks.
@@ 6.2 Language Models
Starting with available open-weight language models, the goal is to adapt them to perform a selected NLU task. Two types of trained models are used: base and instruction-tuned models, where the latter have additionally been trained to follow users’ intent when prompted, such as answering questions or responding to instructions. Specifically, we consider the following language models: gemma-2b and gemma-2b-it from the Gemma family (Team et al., 2024), and mistral-7b and mistral-7b-it from the Mistral [^3] We use v0.3 for both models. family (Jiang et al., 2023). The "it" suffix refers to instruction-tuned models. They are of particular interest for our experiments as they will show results for models with two different numbers of parameters. All models are readily available text-to-text, decoder-only transformer models with open weights, downloadable from Hugging Face.
Note that when using an instruction-following (it) model, inputs are formatted according to the conventions established when training the model on instructions.
@@ 7.1 IMDB
Table 3: Loss and accuracy on the SST-2 task after applying our SpaRTA approach to different types of parameters in the gemma-2b model, given that the same number of trainable parameters are selected. Results are averaged across 10 random seeds.
gemma-2b gemma-2b-it mistral-7b mistral-7b-it method density (%) loss acc. loss acc. loss acc. loss acc. PT (zero-shot) 0.425 92.0% 0.435 86.1% Full FT 100% 0.092 96.9% 0.107 96.3% 0.080 97.4% 0.080 97.3% SpaRTA 5% 0.096 96.7% 0.105 96.3% 0.084 97.2% 0.081 97.5% SpaRTA 0.5% 0.096 96.7% 0.103 96.4% 0.087 96.9% 0.087 97.0% SpaRTA 0.05% 0.106 96.3% 0.114 96.2% 0.080 97.3% 0.076 97.4% LoRA $\approx$0.05% 0.101 96.5% 0.113 96.2% 0.086 97.1% 0.081 97.1%
Table 4: Test loss and accuracy of models adapted to the IMDB dataset with different fine-tuning methods. We also report zero-shot performance of instruction following PT models. For training details, see Table 19.
Model: gemma-2b-it QNLI RTE SST2 QQP MNLI MRPC COLA BoolQ method targets loss acc. loss acc. loss acc. loss acc. loss acc. loss acc. loss mcc loss acc. avg. PT (zero-shot) 1.34 59.5 1.77 56.3 0.75 60.7 1.41 48.3 2.91 30.9 0.91 64.6 0.85 0.7 0.88 68.5 44.7 SpaRTA (5%) ALL 0.17 93.7 0.54 81.0 0.14 95.0 0.26 89.0 0.33 87.2 0.35 85.3 0.39 56.2 0.36 84.4 84.0 SpaRTA (0.5%) 0.16 94.0 0.43 80.7 0.15 94.7 0.24 89.7 0.32 87.8 0.34 86.5 0.45 54.9 0.37 84.7 84.1 SpaRTA (0.037%) 0.18 93.1 0.48 76.9 0.16 94.5 0.26 88.8 0.35 86.4 0.36 84.5 0.45 54.8 0.40 83.6 82.8 LoRA (0.037%) Q,V 0.18 93.4 0.44 78.7 0.14 95.1 0.26 89.1 0.33 87.8 0.33 85.4 0.41 55.0 0.36 84.0 83.6 DoRA (0.037%) 0.18 93.4 0.45 78.0 0.15 95.6 0.26 89.0 0.33 87.8 0.33 85.5 0.41 57.1 0.36 84.4 83.8 SpaRTA (0.037%) 0.19 92.8 0.46 77.6 0.16 94.8 0.26 88.6 0.36 86.3 0.36 84.2 0.45 54.5 0.39 83.4 82.8 SpaRTA (0.037%) O,V 0.17 93.5 0.44 80.7 0.15 94.9 0.26 88.9 0.34 87.2 0.35 85.7 0.42 55.9 0.37 85.0 84.0
Table 5: Test loss and accuracy of gemma-2b-it adapted to GLUE and BoolQ datasets with different fine-tuning methods. Standard errors can be found in Table 15, see Appendix D. For training details see Appendix E.2.
Table 4 shows the results for IMDB, where each model is asked to classify a review as positive or negative. Each model is fine-tuned using the following adaptation methods: (i) Full parameter fine-tuning (Full FT) where all model parameters are optimized; (ii) SpaRTA for different density levels 5%, 0.5%, 0.05%, with the last allowing SpaRTA to have approximately the same number of trainable parameters as LoRA; (iii) LoRA with rank $r=8$, equivalent to about $0.05\%$ of trainable parameters compared to the model full parameter size.